# A One Dimensional (1D) Computational Fluid Dynamics Study of Fontan-Associated Liver Disease (FALD)

### v1

@@ Abstract
Fontan-Associated Liver Disease (FALD) is a disorder arising from hemodynamic changes and venous congestion in the liver. This disease is prominent in patients with hypoplastic left heart syndrome (HLHS). Although HLHS patients typically survive into adulthood, they have reduced cardiac output due to their univentricular physiology (i.e., a Fontan circuit). As a result, they have insufficient blood delivery to the liver. In comparison, patients with double outlet right ventricle (DORV), also having a univentricular circuit, have lower incidence of FALD. In this study, we use a patient-specific, one-dimensional computational fluid dynamics (1D-CFD) model to predict hemodynamics in the liver of an HLHS patient and compare predictions with an age- and size-matched DORV control patient. Additionally, we simulate FALD conditions in the HLHS patient to predict hemodynamic changes across various stages of disease progression. Our results show that the HLHS patient has a higher portal venous pressure compared to the DORV patient. This difference is exacerbated as FALD conditions progress. The wall shear stress (WSS) is also higher than normal for the HLHS patient, suggesting vascular remodeling. WSS decreases slightly under FALD conditions, consistent with the development of portal hypertension. Perfusion analysis gives insight into regions of liver tissue at risk for fibrosis development, showing increasing pressures and reduced flow throughout the liver tissue fed by the portal vein under FALD conditions. Our results provide insight into the specific hemodynamic changes in Fontan circulation that can cause FALD.
@@ 1 Introduction
A common complication is the development of Fontan-associated liver disease (FALD) [7, 8, 9, 10]. Almost Fontan patients will develop FALD to some degree, which could lead to portal hypertension, hepatocellular carcinoma, and eventual liver failure [11]. Due to these complications, patients require liver and/or heart transplants around the age of 20 [12, 13].
[…]
FALD results from an increase in venous pressure in the portal vein caused by an increased central venous pressure (CVP) typical for this group of patients. Despite increased portal vein pressure, the pressure gradient is reduced, thereby reducing blood flow in the liver. The result is chronic systemic venous congestion and reduced cardiac output [20, 21, 22, 23].
FALD progresses through several stages before reaching a final, irreversible stage, where the main complications and breakdown of the circulatory system occur [24]. The initial stages of FALD are characterized by liver congestion and sinusoidal dilation. For most patients, this begins at birth and continues after Fontan surgery [25, 26, 27]. Although many patients remain without symptoms, $53\%$ eventually experience painful hepatomegaly or hepatojugular reflux due to reduced cardiac output [4]. The second stage of FALD involves perisinusoidal fibrosis, regenerative nodules, and hepatocellular necrosis, which occurs 5-10 years after surgery [4]. In most patients, the final stage involves advanced fibrosis with portal vein hypertension, leading to inadequate filtration of the blood within the liver [28, 4]. As FALD progresses, these complications lead to increased stiffening of the vasculature [29].
Currently, the underlying pathology and physiology of FALD are not fully understood. Most FALD studies are based on animal models or clinical data from patient cohorts with other chronic liver diseases [30]. Few studies focus on the pathology of FALD. A review by Cieplucha et al. [31] suggests that patients with HLHS develop FALD due to two factors: 1) permanent reduction in cardiac output and altered hemodynamics due to Fontan reconstruction, and 2) reduced oxygen delivery resulting from the aortic reconstruction. Several clinical studies suggest that passive flow in the Fontan circulation to the pulmonary system is analogous to a cardiovascular system with obstruction in the hepatic venous outflow [30, 32]. In a report by [29], elastographies of Fontan patients revealed that the longer a patient lives with a Fontan circuit, the stiffer their liver vasculature will be. However, it is unknown whether the increase in stiffness is due to increased congestion (i.e., elevated pressures in the liver) or the development of fibrosis.
Computational studies of the Fontan circulation are limited. Most computational studies use three-dimensional computational fluid dynamics (3D-CFD) to investigate velocity patterns and power dissipation [33, 34]. While this technique allows the assessment of complex velocity patterns, it is time-consuming and unrealistic for large-body network models. Moreover, most of these studies focus on the Fontan reconstruction rather than the effects of a Fontan circuit on downstream organs.
An alternative is one-dimensional fluid dynamics (1D-CFD) models. These models efficiently forecast wave propagation within extensive networks and can encompass fluid-solid interactions. For example, the study by [35] uses a 1D arterial and venous network with zero-dimensional heart and organ bed models to study variations in Fontan circulations, contrasting Fontan circuits with fenestration versus hepatic vein exclusion. This study indicates enhanced gut flow but did not consider the potential influence of vortices in bifurcations and reconstructed vessel segments. A recent study by Taylor-LaPole et al. [36] uses a patient-specific 1D-CFD network to predict the effect of aortic reconstruction in an HLHS patient compared to a double outlet right ventricle (DORV) control patient. This study finds that the HLHS patient has hypertensive pressures in the brain and reduced flow to the gut. These studies present promising large vessel networks of the Fontan circulation, but explore neither organ-specific networks nor the potential development of FALD.
Computational studies focusing specifically on liver disease in the Fontan circuit are scarce. The study by Trusty et al. [37] uses computational fluid dynamics simulations to quantify the total cavo-pulmonary connection and cross analyze to data from quantified liver biopsies. Although this helps determine effects within the constructed Fontan conduit downstream of the liver, it does not provide detailed information about the liver vasculature itself. The study by Ishizaki et al. [38] develops a vortex flow map to predict blood movement within the right atrium during the cardiac cycle. The study suggests that reduced vortex flow in the right atrium during the late phase of the Fontan operation is correlated with the development of FALD but does not include liver circulation in the model itself.
To our knowledge, no previous computational studies have analyzed the effect of aortic reconstruction on FALD hemodynamics while incorporating a detailed network of liver vasculature. To investigate the effects of a reconstructed aorta on the development of FALD, we compare an HLHS patient to a control single-ventricle patient with double outlet right ventricle disease (DORV). DORV is another congenital heart defect in which the pulmonary artery and aorta connect to the right ventricle. Patients also receive surgeries construct a Fontan circuit, but no surgical reconstruction of the aorta is performed [39]. To investigate the progression of FALD in a Fontan circuit, we impose increased vascular stiffness and resistance on the patient with HLHS, mimicking the effects of increased CVP. We compare liver perfusion between the patient with HLHS without FALD, the control patient with DORV, and the patients with HLHS and varying degrees of FALD. This is done using a 1D-CFD model predicting pressure, flow, and wall shear stress (WSS) in patients with HLHS and DORV. Our detailed liver network model allows us to investigate pressure changes and blood perfusion within the liver, providing new insight into hemodynamics in the liver of a Fontan circuit, including both arterial and portal vein circulations.
2 Methods
This study uses hemodynamic data from a previous study [36] along with computed tomography (CT) images of the liver to construct an arterial network of both the portal and hepatic vasculature. The liver image is from a representative adult male obtained from an open-access repository. Therefore, vascular dimensions were allometrically scaled as done in [36] to match the specific patients in this study. A 1D CFD model is solved within these networks to predict pressure and flow. The solution is mapped to the liver tissue providing insight into tissue perfusion and distributions of pressure. FALD is simulated in patients with aortic reconstruction by increasing downstream resistance, thus inducing increased CVP.
Figure 1: (a) 3D rendering of the hepatic artery (blue-green) and portal vein (red) vasculatures extracted from the CT. Centerlines extracted from VMTK are shown in black. (b) Inflow waveform (mL/s) for the two patients.
2.1 Data and Image analysis
A CT image of the liver region of an adult male volunteer with tumors in the pancreas is used to construct a 3D representation of the vasculature [40]. We assume that the liver vasculature is normal and is not affected by the tumor. The image can be found in the publicly available 3D-IRCADb-01 database (see https://www.ircad.fr/research/data-sets/liver-segmentation-3d-ircadb-01/). From this image, we extract vascular networks distal to the hepatic artery and the portal vein.
Segmentations of the hepatic artery and portal vein networks are generated using the open-source software 3D Slicer developed by Kitware, Inc. (see http://www.slicer.org) [41, 42]. Using built-in tools from 3D Slicer, including painting, thresholding, cutting, and islanding, a 3D rendering of the arterial and portal vein networks is obtained. The image intensities for the hepatic artery and portal vein segmentations are set to 28.05-255.00 Hounsfeld units. These segmentations are exported as STL files from 3D Slicer and imported to Paraview [43], where they are converted to a VTK polygonal data file for further processing. The liver volume is also segmented. The volume is segmented in the same fashion as the vasculature, with image intensity set to 64.00-255.00 Hounsfeld units. The resulting segmentation is exported as an STL file and opened in Paraview, where it is converted to a volumetric mesh and saved as a csv file storing the $x,y,z$ coordinates of the tissue.
We utilize the Vascular Modeling Toolkit (VMTK) to generate centerlines in 3D renderings of the hepatic artery and portal vein network [44]. This software places maximum inscribed spheres at each $x,y,z$ coordinate of the vessels and approximates the medial axis with a Voronoi diagram [44, 45]. Centerlines are determined as the minimal paths along the inverse of the spheres’ radii [44]. Inlet and outlet boundaries are manually set, and VMTK works recursively from the terminal to the proximal vessel(s), generating centerlines.
Figure 2: Schematic of the Fontan circulation, including the liver flow. The portal vein supplies $70\%$ and the hepatic artery $30\%$ of the flow to the liver.
Labeled Directed Graph We use in-house algorithms in MATLAB to extract network information from VMTK and construct a labeled directed tree with edges defining vessels, nodes, and junctions. Each vessel has labels that denote its radius and length [46, 47]. VMTK defines the points of centerlines as $x$, $y$, and $z$ coordinates along each vessel. The locations at which two central lines interact are junctions. They are adjusted to appear in the ostium’s center using the junction correction algorithm by Bartolo & Taylor-LaPole et al. [48]. The vessel and its labels are integrated into a directed tree, defining a connectivity matrix that connects the vessel to its parent and daughters. Optimal vessel dimensions, i.e., radii and length, are determined using an in-house algorithm utilizing statistical changepoints [48].
Hemodynamics Data Approximately 25% of cardiac output enters the liver [14, 16]. The inflow to the liver is determined by scaling the measured cardiac output [36] to this amount. The DORV patient has a total hepatic inflow of 1.05 L/min and the HLHS patient has an inflow of 1.27 L/min. Of these, 70% enters through the portal vein and 30% through the hepatic artery. Figure 1 shows the scaled inflow waveforms, and Table 1 lists the flows for each patient.
Table 1: Flow and stroke volume entering the hepatic artery and portal vein. All quantities are listed as L/min. For each patient, the cardiac output values are taken from [36].
HLHS DORV
Total cardiac output 5.08 4.06
Portal vein inflow 0.89 0.74
Hepatic artery inflow 0.39 0.31

### v2

@@ Abstract
Fontan-Associated Liver Disease (FALD) is a disorder arising from hemodynamic changes and venous congestion in the liver. This disease is prominent in patients with hypoplastic left heart syndrome (HLHS). Although HLHS patients typically survive into adulthood, they have reduced cardiac output due to their univentricular physiology (i.e., a Fontan circuit). As a result, they have insufficient blood delivery to the liver. In comparison, patients with double outlet right ventricle (DORV), also having a univentricular circuit, have lower incidence of FALD. In this study, we use a patient-specific, one-dimensional computational fluid dynamics (1D-CFD) model to predict hemodynamics in the liver of an HLHS patient and compare predictions with an age- and size-matched single-ventricle Fontan DORV control patient. Additionally, we simulate FALD conditions in the HLHS patient to predict hemodynamic changes across various stages of disease progression. Our results show that the HLHS patient higher hepatic arterial pressure compared to the DORV patient. This difference is exacerbated as FALD conditions progress. HLHS patients also have higher average portal pressures than the DORV. The wall shear stress (WSS) is higher in the hepatic network for the simulated FALD patients. WSS is slightly decreased in the portal network for the HLHS patients, consistent with the development of portal hypertension. Perfusion analysis gives insight into regions of liver tissue at risk for fibrosis development, showing increasing pressures and reduced flow throughout the liver tissue fed by the portal vein under FALD conditions. Our results provide insight into the specific hemodynamic changes in Fontan circulation that can cause FALD.
@@ 1 Introduction
The Fontan procedure is a life-saving surgery for single ventricle patients. Single ventricle disease is a category of rare congenital heart defects that affect about 5 in 100,000 babies [Fixler2010, Reller2008]. The most common type is hypoplastic left heart syndrome (HLHS). HLHS patients are born with an underdeveloped left ventricle and aorta, leaving them with no way to pump oxygenated blood sufficiently to the body [noonan1958hypoplastic]. Patients undergo a series of surgeries from their first weeks of life to age three, resulting in a fully functioning univentricular circulatory system [tellez2018fontan]. This system is called the Fontan circuit. For these patients, the single functioning right ventricle supports the pulmonary and systemic circulations; the ventricle actively pumps blood into the systemic circulation, while the pulmonary circulation is driven by a passive pressure gradient from the systemic veins to the lungs [fontan1971surgicalrepair]. This leads to an elevation in central venous pressure (CVP) [leeshahjehan2023fontancompletion]. Approximately $61\%–85\%$ of patients with Fontan circuits survive until early adulthood, but these patients will suffer long-term complications.
A common complication is the development of Fontan-associated liver disease (FALD) [khairy2008long, downing2017long, schilling2016fontan, kverneland2018five]. Almost all Fontan patients will develop FALD to some degree, which could lead to portal hypertension, hepatocellular carcinoma, and eventual liver failure [emamaullee2020fontan]. Due to these complications, patients require liver and/or heart transplants around the age of 20 [vaikunth2019short, deLange2023].
Blood enters the liver through two conduits and is drained by the hepatic veins, making it a complex system to model. The hepatic artery, branching from the abdominal aorta, transports $25-30\%$ of blood to the liver, while the portal vein supplies the other $70-75\%$ [lautt2009hepatic, eipel2010regulation, rocha2012liver]. The portal vein and hepatic artery systems have a significant pressure difference. The hepatic artery has a mean pressure of $90$ mmHg, and the portal vein has a mean pressure of $10$ mmHg [granger2004circulation]. Studies show that increased pressures of the portal vein or hepatic artery lead to liver damage such as cirrhosis, fibrosis, nonalcoholic fatty liver disease, and portal hypertension [xanthopoulos2019heart, alvarez2011liver]. Understanding the flow and pressure within the liver vasculature due to a Fontan circulation is vital to understanding the pathology of FALD and predicting its development.
FALD results from an increase in the venous pressure in the portal vein caused by an increase in the central venous pressure (CVP) typical of this group of patients. Despite the increased pressure in the portal vein, the pressure gradient is reduced, which reduces blood flow to the liver. The result is chronic systemic venous congestion and reduced cardiac output [navaratnam2016exercise, vandebruaene2015effect, oka2023liver, mazza2021pathophysiology].
FALD progresses through several stages before reaching a final, irreversible stage, where the main complications and the breakdown of the circulatory system occur [gordonwalker2019fontanliver]. The initial stages of FALD are characterized by liver congestion and sinusoidal dilation. For most patients, this begins at birth and continues after Fontan surgery [goldberg2017hepaticfibrosis, johnson2013identifying, schwartz2012hepatic]. Although many patients remain without symptoms, $53\%$ eventually experience painful hepatomegaly or hepatojugular reflux due to reduced cardiac output [tellez2018fontan]. The second stage of FALD involves perisinusoidal fibrosis, regenerative nodules, and hepatocellular necrosis, which occurs 5-10 years after surgery [tellez2018fontan]. In most patients, the final stage involves advanced fibrosis with portal vein hypertension, leading to inadequate filtration of the blood within the liver [rychik2002relation, tellez2018fontan]. As FALD progresses, these complications lead to increased stiffening of the vasculature [Daniels2014].
Currently, the underlying pathology and physiology of FALD are not fully understood. Most FALD studies are based on animal models or clinical data from patient cohorts with other chronic liver diseases [delange2023fontanliverdisease]. Few studies focus on the pathology of FALD. A review by Cieplucha et al. [cieplucha2022fontanliverdisease] suggests that patients with HLHS develop FALD due to two factors: 1) permanent reduction in cardiac output and altered hemodynamics due to Fontan reconstruction, and 2) reduced oxygen delivery resulting from reduced cardiac output. It has been hypothesized that the reconstructed aorta of HLHS Fontan patients further reduces cardiac output compared to other Fontan patients without reconstruction. This is due to abnormal aortic flow leading to power losses and stagnant flow in the aorta [sundareswaran2006] and increased aortic stiffness over time [Walser, Rychik]. Several clinical studies suggest that passive flow in the Fontan circulation to the pulmonary system is analogous to a cardiovascular system with obstruction in the hepatic venous outflow [delange2023fontanliverdisease, schleiger2020evaluation]. In a report by [Daniels2014], elastographies of Fontan patients revealed that the longer a patient lives with a Fontan circuit, the stiffer their liver vasculature will be. However, it is unknown whether the increase in stiffness is due to increased congestion (i.e., elevated pressures in the liver) or the development of fibrosis.
Computational studies of the Fontan circulation are limited. Most computational studies use three-dimensional computational fluid dynamics (3D-CFD) to investigate velocity patterns and power dissipation [marsden2007effectsofexercise, ahmed2021interventionalplanning]. While this technique allows the assessment of complex velocity patterns, it is time-consuming and unrealistic for large-body network models. Moreover, most of these studies focus on the Fontan reconstruction rather than the effects of a Fontan circuit on downstream organs.
An alternative is one-dimensional fluid dynamics (1D-CFD) models. These models efficiently forecast wave propagation within extensive networks and can encompass fluid-solid interactions. For example, the study by [puelz2017computationalstudy] uses a 1D arterial and venous network with zero-dimensional heart and organ bed models to study variations in Fontan circulations, contrasting Fontan circuits with fenestration versus hepatic vein exclusion. This study indicates enhanced gut flow but did not consider the potential influence of vortices in bifurcations and reconstructed vessel segments. A recent study by Taylor-LaPole et al. [alyssa2022] uses a patient-specific 1D-CFD network to predict the effect of aortic reconstruction in two single ventricle Fontan patients: an HLHS patient and a double outlet right ventricle (DORV) patient. The latter, which serves as the control, has the same physiology as the HLHS patient, but does not have a reconstructed aorta. DORV is a congenital heart defect in which the pulmonary artery and aorta connect to the right ventricle; many of these patients also have a nonfunctioning left ventricle [goo2021double]. This study finds that the HLHS patient has hypertensive pressures in the brain and reduced flow to the gut. These studies present promising large vessel networks of the Fontan circulation, but explore neither organ-specific networks nor the potential development of FALD.
Computational studies focusing specifically on liver disease in the Fontan circuit are scarce. The study by Trusty et al. [trusty2018impact] uses computational fluid dynamics simulations to quantify the total cavo-pulmonary connection and cross analyze to data from quantified liver biopsies. Although this helps determine effects within the constructed Fontan conduit downstream of the liver, it does not provide detailed information about the liver vasculature itself. The study by Ishizaki et al. [ishizaki2018prediction] develops a vortex flow map to predict blood movement within the right atrium during the cardiac cycle. The study suggests that reduced vortex flow in the right atrium during the late phase of the Fontan operation is correlated with the development of FALD but does not include liver circulation in the model itself.
To our knowledge, no previous computational studies have analyzed the effect of aortic reconstruction on FALD hemodynamics while incorporating a detailed network of liver vasculature. To investigate the effects of a reconstructed aorta on the development of FALD, we compare the two single ventricle patients with Fontan circulations from [alyssa2022]: an HLHS patient and a DORV patient. Both of these patients have a Fontan circuit, but the DORV patient did not have a surgical reconstruction of the aorta [goo2021double]. Figure 1 shows the two patients’ along with the connectivity of a healthy control with a two-sided heart.
Figure 1: (a) healthy heart with two ventricles. (b) hypoplastic left heart syndrome (HLHS) heart. (c) double outlet right ventricle (DORV) heart with a ventricular septal defect. (d) HLHS heart with a Fontan repair. This patient has undergone aortic reconstruction. (e) DORV heart with a Fontan repair. Patients in (d) and (e) both have a total cavopulmonary connection with the main pulmonary artery connected to the inferior vena cava.
To investigate the progression of FALD in a Fontan circuit, we impose increased vascular stiffness and resistance on the patient with HLHS, mimicking the effects of increased CVP. We compare liver perfusion between the patient with HLHS without FALD, the control patient with DORV, and the patients with HLHS and varying degrees of FALD. This is done using a 1D-CFD model predicting pressure, flow, and wall shear stress (WSS) in patients with HLHS and DORV. Our detailed liver network model allows us to investigate pressure changes and blood perfusion within the liver, providing new insight into hemodynamics in the liver of a Fontan circuit, including both arterial and portal vein circulations. We investigate patient-specific differences by comparing hemodynamic responses with normalized and patient specific cardiac outputs.
2 Methods
This study uses hemodynamic data from a previous study [alyssa2022] along with computed tomography (CT) images of the liver to construct an arterial network of both the portal and hepatic vasculature. The liver image is from a representative adult male obtained from an open-access repository. Therefore, vascular dimensions were allometrically scaled as done in [alyssa2022] to match the specific patients in this study. A 1D CFD model is solved within these networks to predict pressure and flow. The solution is mapped to the liver tissue providing insight into tissue perfusion and distributions of pressure. FALD is simulated in patients with aortic reconstruction by increasing downstream resistance, thus inducing increased CVP.
Figure 2: (a) The Fontan circulation, including flow to the liver. The right heart pumps blood into the aorta, which feed the liver via the portal vein, supplying 70% of the blood, and the hepatic artery, supplying 30% of the blood [rocha2012liver]. From the liver, blood is drained into the systemic veins, which passively is passing through the pulmonary circuit returning to the right heart. (b) 3D rendering of the hepatic artery (blue-green) and portal vein (red) vasculatures extracted from a CT scan from a healthy adult. Centerlines extracted from VMTK are shown in black. The network generated is scaled to the height and weight of the two subjects analyzed here. (c) Inflow waveforms (mL/s) the hepatic artery and portal vein for the two patients. The DORV patient has a greater occurrence of sub-oscillations, that is reflective waves after the initial peak, at the inlet of both networks.
2.1 Data and Image analysis
This study includes simulations of two representative single ventricle subjects with Fontan circulation (Figure 1): a DORV and an HLHS subject. Subjects were paired, matching the outcomes associated with aortic remodeling. Table 1 shows they have similar weight, height, age, blood pressure, and heart rate, but their cardiac output differs. The HLHS patient has a higher cardiac output than the DORV patient.
The DORV and HLHS patients used in this study were part of a larger pilot study using three-dimensional, time-resolved phase contrast cardiac magnetic resonance (CMR) to assess differences in aortic flow properties between functional single right ventricle with and without aortic arch reconstruction. Single ventricle patients who had arch reconstruction (HLHS) and age/sized matched patients with native aortic arches (DORV) with mitral atresia were prospectively recruited. All subjects had a single right ventricle anatomy, less than moderate atrioventricular valve regurgitation, and a total cavopulmonary anastomosis (non-fenestrated). Patients with aortic surgery beyond the initial Norwood and significant collateral burden (<40% "Qp" ), systemic hypertension, and those with heart failure were excluded. Thus, patients, included in this study, had no evidence of failing Fontan physiology by means of cardiac dysfunction or AV valve failure.
A CT image of the liver region of an adult male volunteer with tumors in the pancreas is used to construct a 3D representation of the vasculature [soler2010]. We assume that the liver vasculature is normal and is not affected by the tumor. The image can be found in the publicly available 3D-IRCADb-01 database (see https://www.ircad.fr/research/data-sets/liver-segmentation-3d-ircadb-01/). From this image, we extract vascular networks distal to the hepatic artery and the portal vein.
Segmentations of the hepatic artery and portal vein networks are generated using the open-source software 3D Slicer developed by Kitware, Inc. (see http://www.slicer.org) [Federov12, Kikinis14]. Using built-in tools from 3D Slicer, including painting, thresholding, cutting, and islanding, a 3D rendering of the arterial and portal vein networks is obtained. The image intensities for the hepatic artery and portal vein segmentations are set to 28.05-255.00 Hounsfield units. These segmentations are exported as STL files from 3D Slicer and imported to Paraview [Utkarsh15], where they are converted to a VTK polygonal data file for further processing. The liver volume is also segmented. The volume is segmented in the same fashion as the vasculature, with image intensity set to 64.00-255.00 Hounsfield units. The resulting segmentation is exported as an STL file and opened in Paraview, where it is converted to a volumetric mesh and saved as a csv file storing the $x,y,z$ coordinates of the tissue.
We utilize the Vascular Modeling Toolkit (VMTK) to generate centerlines in 3D renderings of the hepatic artery and portal vein network [Antiga2008]. This software places maximum inscribed spheres at each $x,y,z$ coordinate of the vessels and approximates the medial axis with a Voronoi diagram [Antiga2008, izzo2018vascular]. Centerlines are determined as the minimal paths along the inverse of the spheres’ radii [Antiga2008]. Inlet and outlet boundaries are manually set, and VMTK works recursively from the terminal to the proximal vessel(s), generating centerlines.
Labeled Directed Graph We use in-house algorithms in MATLAB to extract network information from VMTK and construct a labeled directed tree with edges defining vessels, nodes, and junctions. Each vessel has labels that denote its radius and length [Colebank19, Colebank21]. VMTK defines the points of centerlines as $x$, $y$, and $z$ coordinates along each vessel. The locations at which two central lines interact are junctions. They are adjusted to appear in the ostium’s center using the junction correction algorithm by Bartolo & Taylor-LaPole et al. [bartolo2023framework]. The vessel and its labels are integrated into a directed tree, defining a connectivity matrix that connects the vessel to its parent and daughters. Optimal vessel dimensions, i.e., radii and length, are determined using an in-house algorithm utilizing statistical changepoints [bartolo2023framework].
Figure 3: Network adapted from [alyssa2022] used to calculate the inflow to the liver. Flow in vessel 34 feeds the hepatic artery, and flow from vessels 36, 40, and 42 enters the portal vein.
@@ 2.1 Data and Image analysis
Table 1: Patient characteristics, cardiac output and flow to the liver. Patient characteristics and cardiac output are taken from [alyssa2022]. *The cardiac outputs are for the DORV and HLHS patients described in section 2.1. For simulations in section 2.5, the HLHS cardiac output was imposed on the DORV network (simulation 5) and the DORV cardiac output was imposed on the DORV network (simulation 6.
DORV HLHS
Age (years) 12 11
Height (cm) 154.3 151.4
Weight (kg) 59.6 62.0
Blood pressure (mmHg) 110/67 116/65
Cardiac output (L/min)* 4.06 5.08
Total liver CO (L/min)* 1.12 1.30
Portal vein flow (L/min)* 0.71 0.86
Hepatic artery flow (L/min)* 0.41 0.44